# Neural Networks and Friction: Slide, Hold, Learn

### v3

@@ II.3.2 Training parameters
We employ the ADAM optimization method [59] due to its adaptive learning rate capabilities, which helps in converging to the optimal solution more quickly and efficiently than conventional gradient descent methods. A learning rate of 0.001 is chosen as a starting point because it strikes a balance between fast convergence and the risk of overshooting the minimum of the loss function. This value is generally considered a good default and has been proven effective across a wide range of tasks. In order to ensure that each training batch has enough data to represent the dataset’s diversity while keeping the computational demand manageable, we choose to use a batch size of 32. Finally, we use 50 epochs to allow the model opportunity to learn from the data without excessive training time.
@@ II.3.3 Loss function
Inspired by physics-informed neural networks (PINNs) [34], we devise a loss function that reflects physical principles to guide the optimization of the parameters of the network. The conditions that observations impose over the friction coefficient are the direct effect, i.e., logarithmic dependence of velocity magnitude, and healing, i.e., logarithmic increase of the static friction coefficient with time (during holds)
${\partial\mu\over\partial\log V}=\mathrm{c}_{1}\,,\quad{\partial\mu\over\partial\log t}=\mathrm{c}_{2}\,,$ (8)
where $\mu_{s}$ is the static friction coefficient, $c_{1}$ and $c_{2}$ are interface-dependent constants. The latter correspond to $a$ and $(a-b)$ (eq. 3), but we are assuming we do not know that the interface is modeled by these parameters, so shall impose that the change of these relative variations is constant, not that the variations themselves are of a given magnitude. There seems to be a real possibility [21] that these constants may change over time, so, to adapt the framework to this fact, we chose not to enforce the same constants over all the protocols but individually for each sequence, by not mixing information from other different series.
Apart from these two conditions, we must impose that there are no variations until the first jump (as we took for granted that each experiment starts from steady-state conditions) and that the average error is small, in order for the prediction to follow also the general trend in the target data.
@@
Let $\delta\bm{\mu}$ and $\delta\bm{\hat{\mu}}$ represent the vector of normalized predictions and targets (respectively) for a given vector of features (velocity protocol) $\bm{V}$, i.e., $\delta\bm{\mu}=\mathsf{NN}(\bm{V}/V_{ch})$. See that $\delta\bm{\mu},\,\delta\bm{\hat{\mu}},\,\bm{V}\in\mathbb{R}^{N}$ (and $N=100$ in this case). The first part of the loss is the mean-squared error:
@@
where |||||$\cdot$||| | ⋅ | | represents the $L_{2}$ norm of the vector. In overall graph for training, fig. 6, this operation is implemented by the neuron labeled $\mathsf{MS}$.
$\mathsf{loss}_{2}=\left(\delta\mu_{1}-\delta\hat{\mu}_{1}\right)^{2}\,,$ (10)
where the sub-index $1$ refers to the first element of every sequence. The implementation of this operation is depicted in fig. 3: it takes the first sequence value of both targets (Targets) and predictions (Predictions), subtracts each other and squares the difference. Thus, the output of the net is eq. 10.
Figure 3: Scheme of sub-net that implements comparison of initial condition.
The third term of the loss is meant to enforce the direct effect: a sudden peak in friction induced by a jump in leading velocity (which, moreover, is known to be proportional to the logarithm of the ratio of velocities before and after the jump, see section I and appendix A). Assume that there are $N_{J}$ jumps in a given protocol, the term that measures the error between the one predicted by the network and the real one is
@@
the symbol $i$ indexes each jump, while $J_{i}$ refers to the entry of the vector where the jump begins, hence the meaning of the notation $\Delta(\delta\mu)_{J_{i}}=\delta\mu_{J_{i}+1}-\delta\mu_{J_{i}}$. The network depicted in fig. 4 implements this procedure: the inputs are the predictions generated by the GRU net (fig. 2) Predictions, the indices of the entries where velocity jumps begin PosBeforeJ and after PosAfterJ, and the real magnitude of the friction jump $\mathsf{\Delta F_{data}}$, the output is the evaluation of eq. 11.
Figure 4: Sub-net that implements direct effect.
Figure 5: Sub-net that implements logarithmic healing during holds.
@@
This loss term is realized schematically as in fig. 5. The first input is again Predictions, coming from the main network fig. 2, the second one is $\mathsfit{N_{h}}$, which is the position of the entry where the hold ends, it is necessary to extract the predicted value $\delta\mu_{N_{h}}$. The latter is then replicated to form a vector that is subtracted to all predictions. The final input LogTerms weights the healing values as per eq. 12 and “masks” the not-healing ones since it contains zeros that multiply those entries. Once the vector has been weighted, two copies of it are made, shifted to compare successive entries and then summed and squared, yielding as output eq. 13.
@@
$\alpha,\,\beta,\,\gamma\in\mathbb{R}$. It has been found that $\alpha=1$, $\beta=0.1$ and $\gamma=5\cdot 10^{-3}$ yields satisfactory results, yet probably not optimal (see discussion section).
Figure 6: Scheme of the global training network
III Training results
The training was performed in Mathematica [58], using the architecture and hyperparameters discussed in section II and 12000 random sequences corresponding to virtual protocols.
The overall training procedure is depicted in fig. 7 and the final losses are consigned in table 1. The complex loss function eq. 13 was not used at once, but an incremental approach was explored: first, the simplest loss (mean squared error, $\mathsf{ME}$) was used to initialize weights and biases. Second, the network was retrained adding the initial-condition term ($\mathsf{ME+ics}$), which further reduced the loss by 50%. This main part was again extracted and used into a larger scheme that included the direct effect loss term ($\mathsf{ME+ics+de}$). The overall precision suffered, but, as it will be shown, this deterioration of the objective function minimization does not translate into a substantial deterioration of the overall approximation; rather, it seems that this larger value reflects the struggle to capture the jumps’ magnitude. We tried re-training the network from $\mathsf{ME+ics}$ for both direct effect and healing simultaneously ($\mathsf{ME+ics+de+heal}$), to no avail. Apparently, the addition of the healing phenomenology to the loss further deflects from the optimum, but it is believed that this is really a drawback of the RSF model used to generate the synthetic data.
Figure 7: Mean training loss: $\mathsf{ME}$ is trained without initialization, $\mathsf{ME+ics}$ is trained with parameters from $\mathsf{ME}$, $\mathsf{ME+ics+de}$ and $\mathsf{ME+ics+de+heal}$ are trained with parameters from $\mathsf{ME+ics}$.
@@ III Training results
IV Test
Once the training is complete, we take out the main net, which is our ready-to-use $\mathsf{NN}$ function that is to model the interface evolution. We test it using the 3000 sequences that the net did not see during training. For each sequence, the percent error is computed as
$\mathsf{error}(\%)=100{||\mathsf{NN}(\bm{V}/V_{ch})-\delta\bm{\hat{\mu}}/\delta\mu_{ch}||_{2}\over||\delta\bm{\hat{\mu}}/\delta\mu_{ch}||_{2}}\,.$ (15)
The mean test error (averaged over the 3000 sequences) is consigned in table 2 for the four different architectures discussed above. It remains fairly constant, always below 5%, the minimum being attained for the loss combining the averaged square error for the whole vector with the direct enforcement of zero change at $t=0$. Their accuracy can be directly grasped in figs. 10 and 10, which will be commented in the next section.
Finally, fig. 8 compares the error distribution across sequences for both the simplest network and the most complex one. In the latter case, the most current error is $\sim 3\%$ (19% of tests), while $\sim 2\%$ (16% of tests) in the former.
@@ IV Test
Figure 8: Verification error breakdown: percentage of test cases that yield a certain percent error.
V Discussion
As hinted by the test results presented in table 2, figs. 10 and 10 clearly showcase that the GRU-based network $\mathsf{NN}$ is able of closely reproducing the friction dynamics encapsulated in the datasets generated with a customary rate-and-state phenomenological law.
Nevertheless, there are obvious limitations. First, the inset in fig. 10 reveals that the enforcement of the initial condition prevents an initial substantial error, but it does not prevent oscillations even though the velocity remains constant. This issue is deemed of minor importance, since real interface data will certainly come with small levels of noise that the network must filter through. By the way, at this point we will have to consider also the use of regularization to prevent overfitting, a feature that was not necessary as the datasets were noiseless.
The fact that the target (friction coefficient) changes during the steady state when there is no change in the features (loading velocities) is a clear drawback, no matter how slightly the change may be.
Moreover, the absence of the state variable in the dataset brings about inaccuracies in reproducing behavior during holds, despite (or rather, because of) explicitly implementing logarithmic healing in the objective function. Logarithmic healing, eq. 8, entails that the increment of static friction coefficient $\Delta\mu_{s}$ over a timespan $\Delta t$ abides by
@@ V Discussion
where $t$ is the total time elapsed from the arrest. Conversely, if we integrate eq. 2 and then plug it back into eq. 3, a term depending on the initial value of the state variable appears:
@@
which can interfere with the scaling, unless $\theta(t=0)\ll t$. The latter condition may be met in real experiments, in which the holds tend to last for much longer than the characteristic age of the interface, but it is certainly not the case here as the holds are always “short”, i.e., of duration similar to the magnitude of the state variable. This seems to be the most obvious drawback of not considering explicitly the interface state in the data-driven framework. In any case, it does not appear as insurmountable, since (as justified above) in real data the behavior during periods of rest would adjust better to the logarithmic scaling.
Yet another aspect to acknowledge is that the magnitude of direct effect peaks seems to be affected by the holds. See for instance that in fig. 10 there is a relatively short hold period preceded and ended by jumps which are well captured by the nets; conversely, the jumps’ magnitude are significantly underpredicted in fig. 10, which features a longer hold, particularly by the net tuned explicitly for healing.
Some of these issues could be remedied drawing inspiration from PINNs [34] again, by taking advantage of the automatic differentiation [62] to automatically compute gradients of the friction coefficient (output) as a function of the velocities (input) so that they can be directly monitored during training. Not only that, it is logical to expect that automatic differentiation can also enable more elegant approaches to implementing both direct effect and healing in the loss function.
@@
Figure 9: Test example #1: light left axis for loading velocity, dark right axis for friction coefficient evolution. Emphasis on initial conditions. The dashed background line represents the velocity protocol. Protocol with hold playing a secondary role.
Figure 9: Test example #1: light left axis for loading velocity, dark right axis for friction coefficient evolution. Emphasis on initial conditions. The dashed background line represents the velocity protocol. Protocol with hold playing a secondary role.
@@
Figure 10: Test example #2: light left axis for loading velocity, dark right axis for friction coefficient evolution. Emphasis on shortcomings in healing and direct effect. The dashed background line represents the velocity protocol. Protocol with relevant hold.
VI Final Remarks
This study has demonstrated that recurrent neural networks based on Gated Recurrent Units (GRUs) possess the capability to accurately model slide-hold protocols in the context of rate-and-state friction laws, provided that the networks are supplied with adequate data and the loss function is tuned accordingly. Our findings affirm the potential of recurrent neural networks as a powerful tool for understanding complex frictional behavior, showcasing their ability to learn from and predict the dynamics of friction with high fidelity. This success highlights the ability of deep learning approaches to replace customary phenomenological laws for dynamic friction, additionally avoiding the need to define internal variables.
However, our investigation also identified limitations, particularly concerning the handling of logarithmic healing during hold periods. Additionally, we observed a potential conflict between optimizing for healing and for the direct effect within the training process, indicating an area for further refinement in how these fundamental aspects of frictional behavior are integrated into the learning scheme. And let alone that no noise has been included in the training datasets yet, and only protocols made of sudden velocity jumps changes were considered.
This document contains but an exploratory approach to modeling frictional interface via deep learning. Looking forward, there are several promising avenues for enhancing the predictive power and generalizability of recurrent neural networks for dynamic friction. Improvement could be sought through the addition of more nuanced physics-informed loss functions predicated on thermodynamic principles [39, 42] and exploration of alternative or ad hoc[35, 43] RNN architectures). Further parametric studies in terms of network hyperparameters are also warranted, in particular in terms of network depth (stacking GRUs in different ways[35]) and the volumes of available data: unlike cheap synthetic data, experimental data remains relatively scarce; hence it is worth exploring how new techniques [63, 64], aimed at improving data economy in the context of constitutive modeling with neural networks, can help in this case.
Incorporating spring-block dynamics [2] into the model could pave the way for learning from real experimental data, thus circumventing the gap between theoretical models and practical observations. Combined with the use of automatic differentiation [62] (not explored at this point), it could offer a completely new pathway to take advantage of laboratory results to understand friction, replacing the use of experiments to simply fit parameters by producing training data that is the infused in a network that can be used later for predictions.
Moreover, the exploration of advanced machine learning features such as transformers [65] and Long Short-Term Memory (LSTM) networks [66] could offer additional improvements. These technologies, known for their ability to capture long-distance dependencies within data, might provide the key to overcoming some of the current limitations. The promise of deep learning does not stop there: one can foresee that an approach based on neural operators [67] may help also with the problem of generalizing friction from lab conditions to crustal ones.
The end goal is to evolve from traditional phenomenological friction models towards more accurate representations that are likewise derived from the data already used for parameter fitting. This evolution does not entail giving up on a deeper understanding of friction dynamics. Rather, it calls for a complementary integration of solid physical principles with the deep learning paradigm. The rationale for adopting these deep-learning models shall be their demonstrable ability to accurately capture complex interface physics, which, in turn, would justify their use in larger-scale numerical simulations involving those phenomena.
Acknowledgements
[…]
Data availability
Appendix A Alternative loss for direct effect
An alternative loss function to enforce direct effect, inspired by the one used for healing, is also presented. It has not been implemented in the previous analyses, but it is shown for completeness’ sake.
The magnitude of the jump in friction coefficient must be proportional to the logarithm of the ratio of final and initial velocities. Mathematically, we express that as
$\mu(V_{\text{a}})-\mu(V_{\text{b}})\propto$ $\log\left({V_{a}\over V_{b}}\right)$
$\to\delta\mu_{\text{a}}-\delta\mu_{\text{b}}\propto$ $\log\left({V_{a}\over V_{b}}\right)\,,$ (18)
where the sub-indices $a$ and $b$ refer to after and before the velocity jump. This is a physical constraint that does not have to be directly referred to the target values, but it is expressed as a condition that the predictions have to abide by. Thus, for all jumps during a virtual experiment, we should have
$\left[{\delta\mu_{\text{a}}-\delta\mu_{\text{b}}\over\log\left({V_{a}\over V_{b}}\right)}\right]_{J_{1}}=\ldots=\left[{\delta\mu_{\text{a}}-\delta\mu_{\text{b}}\over\log\left({V_{a}\over V_{b}}\right)}\right]_{J_{N_{J}}}\,,$ (19)
where $N_{J}$ refers to the total number of velocity jumps in any given velocity protocol, so $J_{1}$ means “first jump” and $J_{N_{J}}$ “last jump”. Hence, the loss term that promotes the enforcement of the direct effect, expressed in this way, is:
$\mathsf{loss}_{3}=\sum_{i=1}^{N_{J}-1}\left\{\left[{\delta\mu_{\text{a}}-\delta\mu_{\text{b}}\over\log\left({V_{a}\over V_{b}}\right)}\right]_{J_{i+1}}-\left[{\delta\mu_{\text{a}}-\delta\mu_{\text{b}}\over\log\left({V_{a}\over V_{b}}\right)}\right]_{J_{i}}\right\}^{2}\,,$ (20)
in other words, the magnitude of the friction coefficient change over any velocity jump must be the same across jumps.

### v4

@@ II.3.2 Training parameters
We employ the ADAM optimization method [63]. A learning rate of 0.001 is chosen as it is generally considered a good default and has been proven effective across a wide range of tasks. We choose to initially use a batch size of 32 when using most of the data for training, then batch sizes of 10 or just 1 sequence were used, respectively, when training with only 1000 or 100 sequences. Finally, during training we use a maximum of 1000 epochs combined with early stopping (the training is terminated if the loss does not improve after 20 epochs, for larger datasets, or 50 epochs, for smaller ones), combined with gradient clipping to avoid erratic exploration of the loss landscape (maximum magnitude 1), and saving the weights yielding the minimum loss to be the ones used later in inference mode.
@@ II.3.3 Loss function
The aphorism “the Law is a teacher” is often attributed to Saint Augustine of Hippo [64, 65], an early Christian philosopher. In the case of neural networks, “the loss is the teacher”. Inspired by physics-informed neural networks (PINNs) [38], we devise a loss function that takes advantage of automatic differentiation [66] and reflects physical principles to guide the optimization of the parameters of the network.
@@
Let $\delta\bm{\mu}$ and $\delta\bm{\hat{\mu}}$ represent the vector of normalized predictions and targets (respectively) for a given vector of features (velocity protocol) $\bm{V}$, i.e., $\delta\bm{\mu}=\mathsf{NN}(\bm{V}/V_{ch})$. See that $\delta\bm{\mu},\,\delta\bm{\hat{\mu}},\,\bm{V}\in\mathbb{R}^{N}$ (and $N=100$ in this case). The first part of the loss is the mean-absolute error (MAE):
$\mathsf{loss}_{1}={||\delta\bm{\mu}-\delta\bm{\hat{\mu}}||\over N}\,,$ (4)
where |||||$\cdot$||| | ⋅ | | represents the $L_{1}$ norm of the vector. This term leads the parameter adjustment during training to follow the available data as closely as possible. The extra loss terms are aimed to choose among the many ways that there are to minimize this average error.
$\mathsf{loss}_{2}=\left|\delta\mu_{1}-\delta\hat{\mu}_{1}\right|\,,$ (5)
where the sub-index $1$ refers to the first element of every sequence. This term will be useful to stabilize the first predictions, but will be deactivated if there is noise in the target data.
To further stabilize the initial prediction, we also enforce the initial gradient to be zero:
$\mathsf{loss}_{3}={\partial(\delta\mu_{1})\over\partial V}\,,$ (6)
since the velocity should not change in the first step. The conditions that experimental evidence impose over the friction coefficient are the direct effect, i.e., logarithmic dependence of velocity magnitude, and healing, i.e., logarithmic increase of the static friction coefficient with time (during holds)
${\partial\mu\over\partial\log V}=\mathrm{c}_{1}\,,\quad{\partial\mu_{s}\over\partial\log t}=\mathrm{c}_{2}\,,$ (7)
where $\mu_{s}$ is the static friction coefficient, $c_{1}$ and $c_{2}$ are interface-dependent constants. According to eq. 1, the direct effect is controlled by $a$ and healing by $(a-b)$, but we are assuming we do not know that the interface is modeled by these parameters, so shall impose that the change of these relative variations is constant, not that the variations themselves are of a given magnitude.
The third term of the loss is meant to enforce the direct effect: a sudden peak in friction induced by a jump in leading velocity (which, moreover, is known to be proportional to the logarithm of the ratio of velocities before and after the jump, see section I). First, we note that the logarithmic dependence implicit in eq. 7 is equivalent to
$\Delta\left[V{\partial\mu\over\partial V}\right]_{t_{1},t_{2}}=\left(\left[V{\partial\mu\over\partial V}\right]_{t_{1}}-\left[V{\partial\mu\over\partial V}\right]_{t_{2}}\right)=0\,,$ (8)
for any two successive instants $t_{1}$ and $t_{2}$, the derivatives at either time can be efficiently computed using the automatic differentiation capabilities of pyTorch. Assuming that the training data has enough resolution to capture every velocity jump, we can safely choose $t_{1}$ and $t_{2}$ to be any two consecutive points, and we must have that the logarithmic increment to be zero to ensure that the variations of the friction coefficient with respect to velocity happens in logarithmic fashion. Since that should be the case for every two successive data points, the loss term for direct effect must be
$\mathsf{loss}_{4}=\sum_{i=1}^{N-1}\Delta\left[V{\partial(\delta\mu)\over\partial V}\right]_{t_{i},t_{i+1}}\,.$ (9)
The terms corresponding to holds should not enter the previous sum, as there the dominant phenomenon is not the direct effect but healing. Conveniently, the fact that $V\approx 0$ during holds automatically acts as a “mask” [33], effectively removing those terms from the loss.
Finally, let us adress healing. Since none of the two laws feature exactly logarithmic evolution and they assign different values of static friction, it was decided (after some tests, see appendix A) that it works better to leave the network to learn this behavior directly from the minimization of the average error. Future work will be occupied with tailoring the architecture, the training or both to deal with this phenomenon.
$\mathsf{loss}=\lambda_{1}\mathsf{loss}_{1}+\lambda_{2}\mathsf{loss}_{2}+\lambda_{3}\mathsf{loss}_{3}+\lambda_{4}\mathsf{loss}_{4}\,,$ (10)
$\alpha,\,\beta,\,\gamma\in\mathbb{R}$. It has been found that $\lambda_{1}=1$, $\lambda_{2}=\lambda_{3}=0.1$ ($\lambda_{2}=0$ whenever noise is present) and $\lambda_{4}=0.01$ yield satisfactory results. When dealing with noisy data, an extra L2-regularization term is added, with $\lambda_{\text{reg}}=10^{-4}$.
III Training results
For illustration purposes, the overall training procedure is depicted in fig. 2 and the final losses are consigned in table 1 only for the noiseless data. The same loss function was used, but the batch size was changed: 32 for 10000 sequences, 10 for 1000 and 1 for both 100 and 10. This just means that the weights and biases were updated more often the smaller the dataset. For each dataset size, the loss can be reduced to the same final value by tuning this parameter. Similar trends are observed for the noisy data, both for the slip and aging laws. Remarkably, to attain that loss level with noisy data one was to train for many more epochs ($\sim 1000$) and the early-stop parameter (“patience”) has to be increased to 50.
Figure 2: Loss during training: aging and slip laws, for 50 training experiments and 5000.
Table 1: Using the aging law and noiseless data: final loss value and validation error for different training dataset sizes.
Training Dataset Size $10^{1}$ $10^{2}$ $10^{3}$ $10^{4}$
Minimal Loss 0.17 0.11 0.08 0.11
Average Error (%) 4.6 1.2 0.62 0.67
Table 2: Average validation error (%) for different training dataset sizes (test dataset corresponds to sequences generated with the aging law): comparison aging and slip laws (same architecture and training hyperparameters).
Training Dataset Size $10^{1}$ $10^{2}$ $10^{3}$ $10^{4}$
Aging 7.2 1.9 2.2 1.0
Slip 4.42 2.9 3.7 2.5
IV Test
Once the training is complete, we take out the main net, which is our ready-to-use $\mathsf{NN}$ function that is to model the interface evolution. We test it using the 3000 sequences, generated with the aging law eq. 2a, that no net saw during training. For each sequence, the percent error is computed as
$\mathsf{error}(\%)=100{||\mathsf{NN}(\bm{V}/V_{ch})-\delta\bm{\hat{\mu}}/\delta\mu_{ch}||_{2}\over||\delta\bm{\hat{\mu}}/\delta\mu_{ch}||_{2}}\,,$ (11)
the underlying truth is considered to be the noiseless data (this also helps to verify the lack of overfitting).
The mean test error (averaged over the 3000 sequences) is consigned in table 1 for noiseless data (only aging law) and in table 2 for noisy data (both aging and slip laws). For noiseless data, the error is approximately $50\%$ smaller
V Discussion
As hinted by the test results presented in tables 1 and 2, figs. 3, 5 and 5 clearly showcase that the GRU-based networks are able of closely reproducing the friction dynamics encapsulated in the datasets generated with a customary rate-and-state phenomenological law. Remarkably, the aforementioned tables and plots evince that reasonable amounts of training data ($\sim 100$ experiments) are enough for the network to learn the dynamics. This is a promising result, since the next step in this research effort is to learn from laboratory experiments, which are much more limited in number.
Clearly, the loss term associated to logarithmic velocity dependence, eq. 9, is doing its part in concentrating gradients where the sudden changes in velocity magnitude are detected without compromising the transient evolution after velocity steps.
Nevertheless, there still are obvious limitations. First, fig. 3 reveals that the network trained with slip-law data struggles to capture the direct effect right after the hold, no matter how much training data is available. Conversely, the network whose parameters were adjusted with aging-law datasets performs much better, see fig. 5. Their different performance is also illustrated by fig. 5. Therein, for the same amounts of training data (just 100 sequences) the aging-trained model captures better the drop in friction when hold starts and the peak after re-sliding. This behavior could be explained by the fact that the slip-law training data assigns the same value to the static coefficient when the hold starts regardless of the prior dynamic magnitude[19], while this value changes as a function of that magnitude in the case of the aging law. Hence, this may build extra flexibility in the aging-trained net.
The previous discussion highlights the need to develop more tailored loss functions to better account for healing.
Figure 3: Test example #1: light right axis for loading velocity, dark right axis for friction coefficient evolution. Emphasis on capacities of neural network based on slip-law datasets (100, 1000 and 10000 sequences). The dashed background line represents the velocity protocol.
Figure 4: Test example #2: light right axis for loading velocity, dark right axis for friction coefficient evolution. Emphasis on capacities of neural network based on aging-law datasets (100, 1000 and 10000 sequences). The dashed background line represents the velocity protocol.
Figure 5: Test example #3: light right axis for loading velocity, dark right axis for friction coefficient evolution. Emphasis on comparing network trained in slip-law dataset and aging-law dataset for the same amount of training sequences (100). The dashed background line represents the velocity protocol. Protocol with relevant hold.
VI Final Remarks
This study has demonstrated that recurrent neural networks based on Gated Recurrent Units (GRUs) possess the capability to accurately model slide-hold protocols in the context of rate-and-state friction laws, provided that the networks are supplied with adequate data and the loss function is tuned accordingly. These findings affirm the potential of recurrent neural networks as a promising tool for understanding complex frictional behavior, showcasing their ability to learn from and predict the dynamics of friction with high fidelity. This success highlights the ability of deep learning approaches to replace customary phenomenological laws for dynamic friction, additionally avoiding the need to explicitly define internal variables. Another reason to be optimistic is that there are still improvements to be explored when it comes to tailor both the architecture (going beyond “off-the-self” GRU cells) and the training procedure.
However, our investigation also identified limitations. Particularly, we observed a potential conflict between optimizing for healing and for the direct effect within the training process, indicating an area for further refinement in how these fundamental aspects of frictional behavior are integrated into the learning scheme. Moreover, even though I considered both noiseless and noisy data, the levels of synthetic noise may not be representative of the intrinsic variability that one may expect in laboratory settings.
This document contains but an exploratory approach to modeling frictional interface via deep learning. Looking forward, there are several avenues for enhancing the predictive power and generalizability of recurrent neural networks for dynamic friction. Improvement could be sought through the addition of more nuanced physics-informed loss functions predicated on thermodynamic principles [43, 46] and exploration of alternative or ad hoc[39, 47] RNN architectures). Further parametric studies in terms of network hyperparameters are also warranted, in particular in terms of network depth (e.g., stacking GRUs in different ways[39] to make each focus on capturing different phenomena).
Unlike cheap synthetic data, experimental data remains relatively scarce; hence it is worth exploring how new techniques [67, 68], aimed at improving data economy in the context of constitutive modeling with neural networks, can help in this case. It is reassuring for the future steps to know that just dozens of experiments (in lieu of hundreds or thousands of them) may be enough to learn interface dynamics. These volumes of information can be generated by isolated laboratories, without requiring large-scale multi-group collaborations. Notwithsatnding that, we venture the possibility of coordination between experimentalists to generate richer datasets that can confidently be used to train models for particular interface conditions.
Incorporating spring-block dynamics and stick-slip [2] into the model will be another step for learning from real laboratory data beyond gouge experiments, thus considering other type of interfaces as e.g. those in PMMA friction.
Moreover, the exploration of advanced machine learning features such as transformers [69] and Long Short-Term Memory (LSTM) networks [70] could offer additional improvements. These technologies might provide the key to overcoming some of the technical current limitations (e.g., having to work with fixed-length sequences). The promise of deep learning does not stop there. On one hand, universal models of friction based on minimal state cells [47] could allow modeling different material interfaces with a single numerical tool; on the other hand, one can foresee that an approach based on neural operators [71] may help also with discretization dependence, having different data sampling frequencies, and with the problem of generalizing friction from lab conditions to crustal ones.
The end goal is to evolve from traditional phenomenological friction models towards more accurate representations that are likewise derived from the data already used for parameter fitting. This evolution does not entail giving up on a deeper understanding of friction dynamics. Rather, it calls for a complementary integration of solid physical principles with the deep-learning paradigm. The rationale for adopting these deep-learning models shall be their demonstrable ability to accurately capture complex interface physics, which, in turn, would justify their use in larger-scale numerical simulations involving those phenomena.
Acknowledgements
Finally, this paper not only verses on but also benefited from artificial intelligence: GPT-4 has provided both help with and input to the text.
Data availability
Appendix A Intial tests
The simpler preliminary studies were performed in Mathematica [72] only for the aging law, eq. 2a. Mathematica’s automatic differentiation capacbilities are limited, so alternative implementations of the loss had to be explored.
A.1 Architecture
The main net is made up by a single GatedRecurrentLayer, followed by a folded LinearLayer. It is the closest implementation in Mathematica to the architecture used in pyTorch.
A.1.1 Training parameters
We employ the ADAM optimization method [63] with learning rate of 0.001, batch size of 32, and 50 epochs of training (no early stop, no gradient clipping).
A.1.2 Loss function
The first part of the loss is the mean-squared error (MSE):
@@ A.1.2 Loss function
where $||\cdot||_{2}$ represents the $L_{2}$ norm of the vector. In overall graph for training, fig. A.4, this operation is implemented by the neuron labeled $\mathsf{MS}$. This prelimianry study revealed a better performance of MAE, eq. 4.
The second part of the loss represents the difference between the initial values of predictions and targets, eq. 5. The implementation of this operation is depicted in fig. A.1: it takes the first sequence value of both targets (Targets) and predictions (Predictions), subtracts each other and squares the difference.
Figure A.1: Scheme of sub-net that implements comparison of initial condition.
The third term of the loss in this case was meant to enforce the direct effect. Assume that there are $N_{J}$ jumps in a given protocol, the term that measures the error between the one predicted by the network and the real one is
@@
the symbol $i$ indexes each jump, while $J_{i}$ refers to the entry of the vector where the jump begins, hence the meaning of the notation $\Delta(\delta\mu)_{J_{i}}=\delta\mu_{J_{i}+1}-\delta\mu_{J_{i}}$. The network depicted in fig. A.2 implements this procedure: the inputs are the predictions generated by the GRU net, the indices of the entries where velocity jumps begin PosBeforeJ and after PosAfterJ, and the real magnitude of the friction jump $\mathsf{\Delta F_{data}}$, the output is the evaluation of eq. 13. This approach was abandoned in favor of eq. 9, which does no require preprocessing the data and whose implementation is more elegant and way less convoluted.
Figure A.2: Sub-net that implements direct effect.
Figure A.3: Sub-net that implements logarithmic healing during holds.
@@
This loss term is realized schematically as in fig. A.3. The first input is again Predictions, coming from the main network, the second one is $\mathsfit{N_{h}}$, which is the position of the entry where the hold ends, it is necessary to extract the predicted value $\delta\mu_{N_{h}}$. The latter is then replicated to form a vector that is subtracted to all predictions. The final input LogTerms weights the healing values as per eq. 14 and “masks” the not-healing ones since it contains zeros that multiply those entries. Once the vector has been weighted, two copies of it are made, shifted to compare successive entries and then summed and squared, yielding as output eq. 15.
@@
$\alpha,\,\beta,\,\gamma\in\mathbb{R}$. It was found that $\alpha=1$, $\beta=0.1$ and $\gamma=5\cdot 10^{-3}$ yields satisfactory results, which at the time were recognized as suboptimal and subsequently improved by the move to pyTorch.
Figure A.4: Scheme of the global training network
A.2 Training and validation results
The training was performed using 12000 random sequences corresponding to virtual protocols.
The final losses are consigned in table 3.
The complex loss function eq. 15 was not used at once, but an incremental approach was explored: first, the simplest loss (mean squared error, $\mathsf{ME}$) was used to initialize weights and biases. Second, the network was retrained adding the initial-condition term ($\mathsf{ME+ics}$), which further reduced the loss by 50%. This main part was again extracted and used into a larger scheme that included the direct effect loss term ($\mathsf{ME+ics+de}$). The overall precision suffered, but, as it will be shown, this deterioration of the objective function minimization does not translate into a substantial deterioration of the overall approximation; rather, it seems that this larger value reflects the struggle to capture the jumps’ magnitude. Re-training the network from $\mathsf{ME+ics}$ for both direct effect and healing simultaneously ($\mathsf{ME+ics+de+heal}$) was tried, to no avail. The addition of the healing phenomenology to the loss further deflects from the optimum, but this could just be a drawback of the RSF model used to generate the synthetic data.
@@ A.2 Training and validation results
The mean test error (averaged over the 3000 sequences) is consigned in table 4 for the four different architectures discussed above. It remains fairly constant, always below 5%, the minimum being attained for the loss combining the averaged square error for the whole vector with the direct enforcement of zero change at $t=0$. Their accuracy can be directly grasped in figs. 3 and 5, which will be commented in the next section.
Finally, fig. A.5 compares the error distribution across sequences for both the simplest network and the most complex one. In the latter case, the most current error is $\sim 3\%$ (19% of tests), while $\sim 2\%$ (16% of tests) in the former.
@@
Figure A.5: Verification error breakdown: percentage of test cases that yield a certain percent error.
First, the inset in fig. A.7 reveals that the enforcement of the initial condition prevents an initial substantial error, but it does not prevent oscillations even though the velocity remains constant.
Second, the absence of the state variable in the dataset brings about limitations in reproducing behavior during holds, despite (or rather, because of) explicitly implementing logarithmic healing in the objective function. This is evinced in fig. A.7 (see inset). Logarithmic healing, eq. 7, entails that the increment of static friction coefficient $\Delta\mu_{s}$ over a timespan $\Delta t$ abides by
@@
where $t$ is the total time elapsed from the arrest. Conversely, if we integrate eq. 2a and then plug it back into eq. 1, a term depending on the initial value of the state variable appears:
@@
which can interfere with the scaling, unless $\theta(t=0)\ll t$. The latter condition may be met in real experiments, in which the holds tend to last for much longer than the characteristic age of the interface, but it is certainly not the case here as the holds are always “short”, i.e., of duration similar to the magnitude of the state variable.
In any case, this is not an insurmountable drawback, since in laboratory data the behavior during periods of rest would adjust better to the logarithmic scaling.
@@
Figure A.6: Test example #1: light right axis for loading velocity, dark right axis for friction coefficient evolution. Emphasis on initial conditions. The dashed background line represents the velocity protocol. Protocol with hold playing a secondary role.
@@
Figure A.7: Test example #2: light right axis for loading velocity, dark right axis for friction coefficient evolution. Emphasis on shortcomings in healing and direct effect. The dashed background line represents the velocity protocol. Protocol with relevant hold.